OpenStudy (anonymous):

PLEASE HELP<3 I give medals Solve the system: 5x + y = 13 3x = 15 - 3y

OpenStudy (bahrom7893):

you can divide the 2nd equation by 3: x = 5 - y

OpenStudy (lgbasallote):

GO BAHROM7893!!!!

OpenStudy (bahrom7893):

Now plug that value into the first equation: 5*(5-y) + y = 13 25 - 5y + y = 13 -4y = 13-25 -4y = -12 y = 3

OpenStudy (bahrom7893):

x = 5 - y = 5 - 3 = 2

OpenStudy (anonymous):

O.o WEll thanks you guys but idont know what that means. Why are you guys so smart? Would either of you like to marry me lololol? <3

OpenStudy (anonymous):

haha cute proposal :D

OpenStudy (bahrom7893):

lol i think im good haha

OpenStudy (bahrom7893):

fool's the smartest on here anyway.. see he's level 77!

OpenStudy (lgbasallote):

well said bahrom :))

OpenStudy (anonymous):

lol, level is just a very rough estimate of how much this community trust you and how much you active one is in OS.

OpenStudy (anonymous):

Real talk. ha i need me a brain in my life<3 I already love the s h i t out of you guyssss.

OpenStudy (anonymous):

fan time

OpenStudy (lgbasallote):

He even has 188 fans. Obviously many people think so too :))

OpenStudy (anonymous):

Score!

OpenStudy (anonymous):

You guys are so much cooler than me ._. im like level 3

OpenStudy (lgbasallote):

does she get a prize? :)))

OpenStudy (lgbasallote):

we were all once level 3 ilike...we just started out sooner than you :D

OpenStudy (anonymous):

mmmmm, i just want a grilled cheese right meow

OpenStudy (lgbasallote):

the higher you go...the longer it takes to level up...i had been obsessing on this site for a week and my level up is slow @_@

OpenStudy (anonymous):

I use to be really good at school work. I do online classes so it's like i dont even go to school. I cant sink in the lessons and i dont learn. So i have to cheat on sites like these ): Naughty Josie